OpenStudy (jamanorthi):

One in every 6 seeds will sprout. Simulate by using a number cube, and estimate the probability that none of a row of 10 seeds will sprout???

OpenStudy (noseboy908):

Firstly, you need to convert the original given number of successful seeds over the total number into a percentage. Any idea how to do that?

OpenStudy (jamanorthi):

No

OpenStudy (noseboy908):

Alright, no problem. Simply divide it out as it's given: 1 in 6, or 1/6. Please state the decimal that yields.

OpenStudy (jamanorthi):

1/6= 16.6666667

OpenStudy (noseboy908):

While I know what you mean, make sure to put the percentage sign after the number if it's a percentage. Otherwise, it looks like an incorrect division problem. (Also, for this one, I'd advise just leaving it in decimal form) So, next, you are given the fact that you have a row of 10 seeds. So, you need to multiply the percentage, or rather the decimal of the percentage, by the total number of seeds being planted. What number do you get now?

OpenStudy (jamanorthi):

1.6666667

OpenStudy (jamanorthi):

Multiply 16.667% by 10 right?

OpenStudy (noseboy908):

Correct!

OpenStudy (noseboy908):

Now, your question is oddly worded. From what you just found, about 1.6 of the seeds will sprout, but it asks what the chances are that all the seeds will fail. So, as far as I am aware, you can't really calculate that. But hold on, let me get a second opinion.

OpenStudy (jamanorthi):

So how do I do that

OpenStudy (noseboy908):

Honestly, I'm not sure it's possible. I've personally never had to calculate the chance of failure, as seen here. That's why I've asked an intelligent mod I am acquainted with to take a look too.

OpenStudy (jamanorthi):

Ok

OpenStudy (zarkon):

\[P(E)=1-P(\overline{E})\]

OpenStudy (jamanorthi):

What does each letter represent?

OpenStudy (zarkon):

For any event E, the probability of E is equal to one minus the probability of its complement

OpenStudy (zarkon):

or just use the binomial distribution

OpenStudy (jamanorthi):

I am confused

OpenStudy (zarkon):

they want you to simulate...so start rolling dice

OpenStudy (mathmale):

Let's make that a bit more concrete. the way I read the problem, the chances that a given seed will sprout are 1 in 6, and so the chances that it won't sprout are (as Zarkon has pointed out) 1-1/6 = 5/6. There are 10 trials. Thus, the chances that zero seeds will sprout is (5/6)^10. Every seed is independent of every other seed.

OpenStudy (zarkon):

http://www.random.org/dice/?num=10

OpenStudy (mathmale):

P(seed will sprout)=1/6 P(seed will not sprout)=5/6 P(2 seeds in a row will not sprout) = (5/6)^2, or (5/6)(5/6). ... ... P(10 seeds in a row will not sprount) = P(zero out of 10 will sprout) = (5/6)^10.

OpenStudy (jamanorthi):

I get it now

OpenStudy (mathmale):

I'm so glad. This probability stuff used to be almost a complete mystery for me, but years of practice took care of that. More power to you!

OpenStudy (jamanorthi):

Thanks :)

OpenStudy (mathmale):

:) with apologies to @Zarkon for having butted in, as I know I did.

OpenStudy (zarkon):

np

OpenStudy (mathmale):

Good sport!